# Kernel Spectral Joint Embeddings for High-Dimensional Noisy Datasets using Duo-Landmark Integral Operators

### v2

@@ A Additional numerical results
For the numerical simulation setup considered in our main text, where in each case the two datasets contain some shared structures, we demonstrate the use of our proposed alignability screening procedure. In particular, throughout this paper, we use $k=30$ for the nearest neighbor parameter in the alignability screening step. For both the simultaneous clustering problem and the nonlinear manifold learning problem, in almost all the settings ($>99\%$), the simulated data pairs passed our alignability screening. Figure F contains the boxplots of the median KNN purity scores within each simulation setting. Our results indicate that the proposed alignability screening does not have a negative effect on the integration of datasets containing (partially) shared structures.
@@
When the pair of datasets do not contain overlapping signals, integrating these datasets using our proposed duo-landmark joint embedding procedure may introduce unwanted distortions to the datasets, creating artificial alignment patterns that may lead to misinterpretations and false conclusions. To demonstrate this potential risk and the power of the proposed alignability screening procedure, we generate a pair of negative control datasets, each containing a distinct manifold structure and there is no (measurable) overlap between the two manifolds. Specifically, the first dataset contains $n_{1}=3000$ samples, uniformly drawn from a “Klein bottle” manifold in 4-dimensional space; see Figure G for an illustration. The second dataset contains $n_{2}=2000$ samples in $\mathbb{R}^{4}$, whose first coordinates are uniformly drawn from $[-1,1]$, and the other coordinates are set as 0. If we directly apply our duo-landmark kernel embedding, we observe an artifical alignment between the two datasets (Figure H); in contrast, the joint embeddings based on the j-kpca do not display similar artificial alignment patterns (Figure I). Now we apply our proposed alignability screening procedure (see Algorithm 1, Step 1) and compute the KNN purity score for each pair of generated data, finding that in all cases (across 100 rounds of simulations), our alignability screening procedure can consistently detect the non-alignability between the two datasets. This suggests that in practice this procedure can effectively avoid generating misleading joint embedding of the two datasets.
@@ D.1 Construction of auxiliary RKHSs
For notional simplicity, we denote these two RKHSs as ̊$\mathring{\mathcal{H}}_{1}$ and ̊$\mathring{\mathcal{H}}_{2}.$
@@
Under the assumption that $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are compactly embedded into some common space $\mathcal{N},$ we can use the conventions in Section D.1 to conclude that these exist some common measurable space $\mathring{\mathcal{N}}$ so that ̊$\mathring{\mathcal{H}}_{1}$ and ̊$\mathring{\mathcal{H}}_{2}$ are compactly embedded into $\mathcal{L}_{2}(\mathring{\mathcal{N}},\mathring{\mathsf{P}}_{1})$ and $\mathcal{L}_{2}(\mathring{\mathcal{N}},\mathring{\mathsf{P}}_{2}),$ respectively. Suppose $\phi(\cdot)$ is an eigenfunction of $\mathcal{K}_{1}$ associated with some nonzero eigenvalue $\lambda$. That is,
@@
By Lemma C.3 and the conventions of Section D.1, we can expand $\phi(\bm{x})$ on the common space $\mathring{\mathcal{N}}$ via $\mathcal{L}_{2}(\mathring{\mathcal{N}},\mathring{\mathsf{P}}_{2})$ using the basis of ̊$\mathring{\mathcal{H}}_{2}.$ Consequently, we can write that for some constants $b_{i}\equiv b_{i}(\phi),\ i\in\mathbb{N}$ and $(b_{i})\in\ell_{2}(\mathbb{R})$ so that
@@ E.1 Hyperparameter selection, computational complexity, and noise detection
For the $k$NN hyperparameter $k$ in the alignability screening step of Algorithm 1, we recommend choosing $k=30$ in practice. This default value was used in all our numerical studies. In our simulation, by comparing the results under different $k\in\{20,25,30,35,40\}$, we found the algorithm’s performance was not sensitive to the choice of this parameter.
@@
Nevertheless, the focus of our current paper is to develop statistically efficient and practically interpretable algorithms for constructing joint embeddings of noisy and nonlinear datasets. To improve numerical efficiency, two potential approaches may be considered. First, one could incorporate randomized numerical linear algebra techniques. However, this presents theoretical challenges. To the best of our knowledge, most existing algorithms and their convergence guarantees assume that the target data matrix is clean and noise-free. To rigorously integrate these techniques into our framework, one would need to analyze their behavior on noisy matrices and establish convergence results that depend on the condition number of the underlying clean matrix and the signal-to-noise ratio. Second, one could consider using non-smooth kernel functions, such as the $k$NN or 001111 kernel, to construct sparse kernel matrices. However, to our knowledge, the convergence properties of such sparse kernel matrices in the context of high-dimensional, noisy, non-smooth manifolds remain unknown. Moreover, the techniques developed in the current paper do not extend to such settings.

### v3

@@ A Additional numerical results
For the numerical simulation setup considered in our main text, where in each case the two datasets contain some shared structures, we demonstrate the use of our proposed alignability screening procedure. In particular, throughout this paper, we use $k=30$ for the nearest neighbor parameter in the alignability screening step. For both the simultaneous clustering problem and the nonlinear manifold learning problem, in almost all the settings ($99\%$), the simulated data pairs passed our alignability screening. Figure F contains the boxplots of the median KNN purity scores within each simulation setting. Our results indicate that the proposed alignability screening does not have a negative effect on the integration of datasets containing (partially) shared structures.
@@
When the pair of datasets do not contain overlapping signals, integrating these datasets using our proposed duo-landmark joint embedding procedure may introduce unwanted distortions to the datasets, creating artificial alignment patterns that may lead to misinterpretations and false conclusions. To demonstrate this potential risk and the power of the proposed alignability screening procedure, we generate a pair of negative control datasets, each containing a distinct manifold structure and there is no (measurable) overlap between the two manifolds. Specifically, the first dataset contains $n_{1}=3000$ samples, uniformly drawn from a “Klein bottle” manifold in 4-dimensional space; see Figure G for an illustration. The second dataset contains $n_{2}=2000$ samples in $\mathbb{R}^{4}$, whose first coordinates are uniformly drawn from [1,1][-1,1], and the other coordinates are set as 0. If we directly apply our duo-landmark kernel embedding, we observe an artifical alignment between the two datasets (Figure H); in contrast, the joint embeddings based on the j-kpca do not display similar artificial alignment patterns (Figure I). Now we apply our proposed alignability screening procedure (see Algorithm 1, Step 1) and compute the KNN purity score for each pair of generated data, finding that in all cases (across 100 rounds of simulations), our alignability screening procedure can consistently detect the non-alignability between the two datasets. This suggests that in practice this procedure can effectively avoid generating misleading joint embedding of the two datasets.
@@ D.1 Construction of auxiliary RKHSs
For notional simplicity, we denote these two RKHSs as $\mathring{\mathcal{H}}_{1}$ and $\mathring{\mathcal{H}}_{2}.$
@@
Under the assumption that $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are compactly embedded into some common space $\mathcal{N},$ we can use the conventions in Section D.1 to conclude that these exist some common measurable space $\mathring{\mathcal{N}}$ so that $\mathring{\mathcal{H}}_{1}$ and $\mathring{\mathcal{H}}_{2}$ are compactly embedded into $\mathcal{L}_{2}(\mathring{\mathcal{N}},\mathring{\mathsf{P}}_{1})$ and $\mathcal{L}_{2}(\mathring{\mathcal{N}},\mathring{\mathsf{P}}_{2}),$ respectively. Suppose $\phi(\cdot)$ is an eigenfunction of $\mathcal{K}_{1}$ associated with some nonzero eigenvalue $\lambda$. That is,
@@
By Lemma C.3 and the conventions of Section D.1, we can expand $\phi(\bm{x})$ on the common space $\mathring{\mathcal{N}}$ via $\mathcal{L}_{2}(\mathring{\mathcal{N}},\mathring{\mathsf{P}}_{2})$ using the basis of $\mathring{\mathcal{H}}_{2}.$ Consequently, we can write that for some constants $b_{i}\equiv b_{i}(\phi),\ i\in\mathbb{N}$ and $(b_{i})\in\ell_{2}(\mathbb{R})$ so that
@@ E.1 Hyperparameter selection, computational complexity, and noise detection
For the kkNN hyperparameter $k$ in the alignability screening step of Algorithm 1, we recommend choosing $k=30$ in practice. This default value was used in all our numerical studies. In our simulation, by comparing the results under different $k\in\{20,25,30,35,40\}$, we found the algorithm’s performance was not sensitive to the choice of this parameter.
@@
Nevertheless, the focus of our current paper is to develop statistically efficient and practically interpretable algorithms for constructing joint embeddings of noisy and nonlinear datasets. To improve numerical efficiency, two potential approaches may be considered. First, one could incorporate randomized numerical linear algebra techniques. However, this presents theoretical challenges. To the best of our knowledge, most existing algorithms and their convergence guarantees assume that the target data matrix is clean and noise-free. To rigorously integrate these techniques into our framework, one would need to analyze their behavior on noisy matrices and establish convergence results that depend on the condition number of the underlying clean matrix and the signal-to-noise ratio. Second, one could consider using non-smooth kernel functions, such as the kkNN or 011 kernel, to construct sparse kernel matrices. However, to our knowledge, the convergence properties of such sparse kernel matrices in the context of high-dimensional, noisy, non-smooth manifolds remain unknown. Moreover, the techniques developed in the current paper do not extend to such settings.